OpenStudy (anonymous):

Solve if possible 3 |x+4|=18

OpenStudy (cwrw238):

|x+4| = 6 this is same as x + 4 = 6 or x + 4 = -6 x = 2 or x = -10

OpenStudy (anonymous):

what cwrw238 said